# The Cohen–Lenstra moments over function fields via the stable homology of non-splitting Hurwitz spaces

### v1

@@ 1.6. Conjectures on stable homology of Hurwitz spaces
The special case of the above conjecture where $G=S_{d}$, $k=1$, and $c_{1}$ is the conjugacy class of transpositions is given in [EVW16, Conjecture 1.5]. The above conjecture was suggested to us by Melanie Wood, who in turn pointed out to us that a statement very closely related to this conjecture was formulated in [EVW12, Corollary 5.8].
@@
Remark 1.6.0.
The authors are currently working on proving additional cases of § 1.6.
@@ 1.9. Acknowledgements
We thank Jordan Ellenberg and Craig Westerland for sharing a number of the ideas they tried while working on this problem. The ideas in paper build heavily upon ideas in the retracted preprint [EVW12], and we were able to draw much inspiration from the ideas presented there. We thank Melanie Wood for suggesting § 1.6, listening to a detailed explanation of the argument, and for helpful comments on the paper. We thank Andrea Bianchi for an extremely helpful suggestion which allowed us to remove an extraneous condition we had on Theorem 4.2.1. We thank Yuan Liu for explaining to us the correct version of the Cohen–Lenstra heuristics in the function field case, and for help understanding the non-abelian Cohen–Lenstra heuristics with roots of unity. We also thank Samir Canning, Kevin Chang, Ben Knudsen, Jef Laga, Jeremy Miller, Sun Woo Park, Anand Patel, Sam Payne, Dan Petersen, Eric Rains, Oscar Randal-Williams, Will Sawin, Mark Shusterman, Nathalie Wahl, and Angelina Zheng for helpful conversations. Landesman was supported by the National Science Foundation under Award No. DMS 2102955. Levy was supported by the NSF Graduate Research Fellowship under Grant No. 1745302, and by the Clay Research Fellowship.
@@ 2.1. The definition of Hurwitz spaces
Let $G$ be a finite group, $c\subset G$ be a union of conjugacy classes, and $B$ be a scheme with $|G|$ invertible on $B$. Let $[\operatorname{Hur}^{G,c}_{n,B}/G]$ denote the Hurwitz stack whose $T$ points $[\operatorname{Hur}^{G,c}_{n,B}/G](T)$ is the groupoid
@@
The definition of Hurwitz spaces we give here is a special case of the definition in [EL23, Definition 2.4.2]. See [EL23, Remark 2.4.3] for an explanation of why these Hurwitz stacks are indeed stacks. It essentially follows from [ACV03, §1.3.2 and Appendix B].
@@
where $D,i,X,$ and $h$ satisfy the properties listed in § 2.1. We also assume the order of inertia of $h$ along $\infty$ is $w$ and define ~$\widetilde{\infty}_{T}$ to be the base change of the section $\widetilde{\infty}$ defined above to $T$. We additionally impose the condition that $h^{\prime}$ is a finite locally free $G$-cover, étale over $\widetilde{\infty}$, such that the composition of $h^{\prime}:X\to\mathscr{P}^{w}_{T}$ with the coarse space map $\mathscr{P}^{w}_{T}\to\mathbb{P}^{1}_{T}$ is $h$, and $t:T\to X\times_{h^{\prime},\mathscr{P}^{w}_{T},\widetilde{\infty}_{T}}T$ is a section of $h^{\prime}$ over $\widetilde{\infty}$.
@@ 4.4. Decomposition of the comparison map
If $c^{\prime}$ is empty, then we claim that both the source and target of the map (4.1) are 00. We claim that some power of $U$ acts by 00 on the source and target, since a large enough power of $U$, $U^{j}$, is in the two sided ideal of $\pi_{*}R$ generated by $\alpha_{\beta}^{\operatorname{ord}(\beta)n_{\beta}}$. This is because by the pigeonhole principle, one can see that any element of $\pi_{0}\operatorname{Hur}^{G,c}_{n}$, written as a product of elements in $c$ for $n$ large enough contains $\max_{\beta\in c}(\operatorname{ord}(\beta)n_{\beta})$ many copies of a single $\beta\in c$. By pulling these to the left using the braid group action, we see the element is divisible by $\alpha_{\beta}^{\operatorname{ord}_{\beta}n_{\beta}}$. Since $U^{j}$ is in the ideal generated by the smallest powers of $\alpha_{\beta}^{\operatorname{ord}(\beta)}$ that act trivially, and hence it acts trivially. Since $U$ acts both as an isomorphism and nilpotently, the source and target of (4.1) with $c^{\prime}=\emptyset$ are both 00.
@@ 4.5. Showing the map is an equivalence
In this section we show that the maps in § 4.4 are an equivalence for each subrack $c^{\prime}\subset c$. In the case that $c^{\prime}=c$, the fact that the map in § 4.4 is an equivalence comes from the fact that the homology of the group completion of $\operatorname{Hur}^{c}$ can be identified with the rack homology of the rack $c$, which is computed in [EG03]. We verify this in § 4.5 For $c^{\prime}\subset c$. In general, we show that completion at $\alpha_{c-c^{\prime}}$ (see § 4.4) corresponds to projecting in pointed spaces to the components of $\operatorname{Hur}^{c^{\prime}}$ at the level of homology, allowing us to again apply § 4.5 for the rack $c^{\prime}$ in this case. Identifying the completion in terms of $\operatorname{Hur}^{c^{\prime}}$ is at the core of our proof, and is proven using a homological epimorphism argument whose main input is a topological one identifying two bar constructions in pointed spaces as homotopy equivalent.
@@
§ 4.5 was essentially already known, and can be found in [RW20, Corollary 5.4]. Although this was stated for conjugacy classes of groups, the proof works equally well for racks. We now spell out the details of this argument for general racks.
@@
By the group completion theorem, $v[\alpha_{c}^{-1}]$ can be identified with the map $\Omega B\operatorname{Hur}^{c}\to\Omega B\operatorname{Conf}^{c}$ on homology. It is then enough to show that the map $\Omega B\operatorname{Hur}^{c}\to\Omega B\operatorname{Conf}^{c}$ is a rational homology equivalence. Since the map $\pi_{1}B\operatorname{Hur}^{c}\to\pi_{1}B\operatorname{Conf}^{c}$ is an equivalence by construction of $\operatorname{Conf}^{c}$, by the Serre spectral sequence of the path loop fibration, it suffices to check $B\operatorname{Hur}^{c}\to B\operatorname{Conf}^{c}$ is a rational homology equivalence.
We will first show that the map
@@
is an equivalence. Note that $\pi_{1}(\psi):\pi_{1}(\Omega B\operatorname{Conf}^{c})\to\pi_{1}(\Omega B\operatorname{Conf}\times_{\Omega B\pi_{0}\operatorname{Conf}}\Omega B\pi_{0}\operatorname{Conf}^{c})$ is an equivalence. Therefore, in order to show $\psi$ is an equivalence, it suffices to show $\psi$ is an equivalence on homology. By the group completion theorem,
@@
Hence, $\psi$ is an equivalence.
@@
It remains then to show that the map $B\operatorname{Hur}^{c}\to B\operatorname{Conf}$ is an equivalence on rational homology. This follows as in [RW20, Corollary 5.4]: This map can be modeled as the map of rack spaces as in [FRS95], from the rack $c$ to the trivial rack *, and the induced map on homology is the map on rack homology induced from this map of racks. It is evident that the map on rack homology is a split surjection (since the associated map of braided vectors spaces has a splitting sending $1$ to $\frac{1}{|c|}\sum_{g\in c}g$), and on the other hand [EG03, Theorem 4.2] shows that the rack cohomology of both sides has rank $1$ in each degree (using that $c$ is connected). Thus the map is an isomorphism. ∎
@@
For $c$ a rack and $c^{\prime}\subset c$ a subset, we define the normalizer of $c$ in $c^{\prime}$ to be $N(c^{\prime}):=\{x\in c|x\triangleright c^{\prime}=c^{\prime}\}$. We say that a subrack $c^{\prime}\subset c$ is self-normalizing if $c^{\prime}=N(c^{\prime})$.
@@
So, it is enough to verify the map (4.7) is a homology equivalence, which follows from § 4.5 below, using that $c^{\prime}$ is self-normalization so $N(c^{\prime})=c^{\prime}$. ∎
@@
Let $c$ be a rack, $c^{\prime}\subset c$ be a subrack with normalizer $N(c^{\prime})$, and $X_{+}$ be $\pi_{0}\operatorname{Hur}^{c^{\prime}}[\alpha_{c^{\prime}}^{-1}]_{+}$, viewed as a $\operatorname{Hur}^{c^{\prime}}_{+}$-bimodule. Then the map
(4.8) $X_{+}\otimes_{\operatorname{Hur}^{c}_{+}}X_{+}\to X_{+}\otimes_{\operatorname{Hur}^{N(c^{\prime})}_{+}}X_{+}$
@@
The map of interest $X_{+}\otimes_{\operatorname{Hur}^{c}_{+}}X_{+}\to X_{+}\otimes_{\operatorname{Hur}^{N(c^{\prime})}_{+}}X_{+}$ has a section induced from the inclusion of racks $N(c^{\prime})\to c$. It thus suffices to show that this section induces a homology equivalence.
We use the notation $\overline{Q}^{*}_{\epsilon}[M,\operatorname{Hur}_{+}^{N(c^{\prime})},N]$, as defined in § A.4 and Theorem A.4.1, which we describe below. By Theorem A.4.1, in order to prove this section is a homology equivalence it suffices to show the inclusion
$\iota_{\epsilon}:\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}_{+}^{N(c^{\prime})},X_{+}]\to\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}_{+}^{c},X_{+}]$
@@
The space $\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{N(c^{\prime})}_{+},X_{+}]$ is defined similarly, except the the labelling function $l$ must land in $N(c^{\prime})\subset c$.
At this point, it may be useful to refer to the notion of two spaces being ind-weakly equivalent, defined in § A.4. Let $S^{c,c^{\prime}}_{\epsilon}$ denote the quotient of the inclusion $\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{N(c^{\prime})}_{+},X_{+}]\to\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$.
@@
Note that any point in $S^{c,c^{\prime}}_{\epsilon}$ that is not the base point can be represented by a point in $\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$ where some $l(p_{i})$ is not in $N(c^{\prime})$.
It will be easier to prove the desired homology equivalence at the level of the associated graded of a certain filtration, which we define next. We define the filtration $F_{\bullet}S^{c,c^{\prime}}_{\epsilon}$ on $S^{c,c^{\prime}}_{\epsilon}$, where, for $j\geq 0$, $F_{j}S^{c,c^{\prime}}_{\epsilon}$ is the subset of $S^{c,c^{\prime}}_{\epsilon}$ consisting of those points where one of $l(p_{k-j}),\ldots,l(p_{k})$ lies in $N(c^{\prime})-c$, together with the base point. We set $F_{-1}S^{c,c^{\prime}}_{\epsilon}$ to be the basepoint, and use $G_{j}S^{c,c^{\prime}}_{\epsilon}$ to denote the associated graded of the filtration $F_{j}S^{c,c^{\prime}}_{\epsilon}/F_{j-1}S^{c,c^{\prime}}_{\epsilon}$, obtained by identifying $F_{j-1}S^{c,c^{\prime}}_{\epsilon}$ with the base point.
@@
We may choose a function $\xi:c-N(c^{\prime})\to c^{\prime}$ such that $g\triangleright\xi(g)\notin c^{\prime}$ for each $g\in c-N(c^{\prime})$, since $g\triangleright c^{\prime}\neq c^{\prime}$ from the definition of the normalizer of $c^{\prime}$.
@@
Suppose $y$ as above lies in $F_{j}Q_{\epsilon}^{c,c^{\prime}}$. Define $v$ to be the largest number so that $l(p_{v})\notin N(c^{\prime})$. By definition of the filtration $F_{j}Q_{\epsilon}^{c,c^{\prime}}$, we have $v\geq k-j$. Let $\epsilon^{\prime}:=\frac{\epsilon}{2}$. We now construct a continuous homotopy $H:F_{j}Q^{c,c^{\prime}}_{\epsilon}\times I\to G_{j}S^{c,c^{\prime}}_{\epsilon^{\prime}}$ as follows: If either $m,n$ is the base point, the point $y$ is equivalent to the base point under relation $(iii)$, and we choose the constant homotopy at the base point. So now assume this is not the case. If $v>k-j$, then the point is in $F_{j-1}Q_{\epsilon}^{c,c^{\prime}}$, so goes to the base point in the associated graded $G_{j}Q_{\epsilon}^{c,c^{\prime}}$. Thus, we may choose the constant homotopy on such points as well. So now we may assume as well that $v=k-j$.

### v2

@@ 1.6. Conjectures on stable homology of Hurwitz spaces
The special case of the above conjecture where $G=S_{d}$, $k=1$, and $c_{1}$ is the conjugacy class of transpositions is a conjecture of Ellenberg-Venkatesh-Westerland [EVW16, Conjecture 1.5]. Even in this special case, our conjecture is slightly stronger because Ellenberg-Venkatesh-Westerland do not conjecture that the stability occurs in a linear range. The above conjecture was suggested to us by Melanie Wood, who in turn pointed out to us that a statement very closely related to this conjecture was formulated in [EVW12, Corollary 5.8].
@@ 1.9. Acknowledgements
We thank Jordan Ellenberg and Craig Westerland for sharing a number of the ideas they tried while working on this problem. The ideas in paper build heavily upon ideas in the retracted preprint [EVW12], and we were able to draw much inspiration from the ideas presented there. We thank Melanie Wood for suggesting § 1.6, listening to a detailed explanation of the argument, and for helpful comments on the paper. We thank Andrea Bianchi for an extremely helpful suggestion which allowed us to remove an extraneous condition we had on Theorem 4.2.1. We thank Yuan Liu for explaining to us the correct version of the Cohen–Lenstra heuristics in the function field case, and for help understanding the non-abelian Cohen–Lenstra heuristics with roots of unity. We also thank Samir Canning, Kevin Chang, Ben Knudsen, Jef Laga, Jeremy Miller, Sun Woo Park, Anand Patel, Sam Payne, Dan Petersen, Eric Rains, Oscar Randal-Williams, Will Sawin, Mark Shusterman, Agata Smoktunowicz, Nathalie Wahl, and Angelina Zheng for helpful conversations. Landesman was supported by the National Science Foundation under Award No. DMS 2102955. Levy was supported by the NSF Graduate Research Fellowship under Grant No. 1745302, and by the Clay Research Fellowship.
@@ 2.1. The definition of Hurwitz spaces
Let $G$ be a finite group, $c\subset G$ be a union of conjugacy classes. Let $B$ be a scheme with $|G|$ invertible on $B$. Assume either that $c$ is closed under invertible powering ($g\in c\implies g^{t}\in c$ for any $t$ prime to $|G|$,) or that $B$ is henselian with residue field $\operatorname{Spec}\mathbb{F}_{q}$ and $c$ is closed under under $q$th powering ($g\in c\implies g^{q}\in c$). Let $[\operatorname{Hur}^{G,c}_{n,B}/G]$ denote the Hurwitz stack whose $T$ points $[\operatorname{Hur}^{G,c}_{n,B}/G](T)$ is the groupoid
@@
When $c=G-\operatorname{id}$, the definition of Hurwitz spaces we give here is a special case of the definition in [EL23, Definition 2.4.2]. See [EL23, Remark 2.4.3] for an explanation of why these Hurwitz stacks are indeed stacks. It essentially follows from [ACV03, §1.3.2 and Appendix B].
The above paragraph shows that $[\operatorname{Hur}^{G,G-\operatorname{id}}_{n,B}/G]$ is an algebraic stack, but does not yet justify the existence of the stacks $[\operatorname{Hur}^{G,c}_{n,B}/G]$ under the suitable powering hypotheses, depending on $B$. First, if $|G|$ is invertible on $B$, and $c$ is closed under invertible powering, then $[\operatorname{Hur}^{G,c}_{n,B}/G]$ is a union of components of the Hurwitz space $[\operatorname{Hur}^{G,G-\operatorname{id}}_{n,B}/G]$ by [LWZB24, Corollary 12.2]. Second, we explain why $[\operatorname{Hur}^{G,c}_{n,B}/G]$ is a union of components of the Hurwitz space $[\operatorname{Hur}^{G,G-\operatorname{id}}_{n,B}/G]$ when $B$ is a Henselian dvr with finite residue field. We can first reduce to the case that $B=\mathbb{F}_{q}$ using that $B$ is Henselian and $[\operatorname{Hur}^{G,G-\operatorname{id}}_{n,B}/G]$ has a smooth proper compactification by [EL23, Corollary B.1.4]. Since $c$ is closed under $q$th powering, the result then holds using by [LWZB24, Theorem 12.1(2)].
@@
where $D,i,X,$ and $h$ satisfy the properties listed in § 2.1. We also assume the order of inertia of $h$ along $\infty$ is $w$ and define ~$\widetilde{\infty}_{T}$ to be the base change of the section $\widetilde{\infty}$ defined above to $T$. We additionally impose the condition that $h^{\prime}$ is a finite locally free $G$-cover, étale over ~$\widetilde{\infty}_{T}$, such that the composition of $h^{\prime}:X\to\mathscr{P}^{w}_{T}$ with the coarse space map $\mathscr{P}^{w}_{T}\to\mathbb{P}^{1}_{T}$ is $h$, and $t:T\to X\times_{h^{\prime},\mathscr{P}^{w}_{T},\widetilde{\infty}_{T}}T$ is a section of $h^{\prime}$ over ~$\widetilde{\infty}_{T}$.
@@ 4.1. Background on racks
Remark 4.1.0.
We may view the map $\triangleright:c\times c\to c$ as a map $\phi:c\to\mathrm{Hom}(c,c),x\mapsto\phi_{x}$. In this setting, § 4.1(2) is equivalent to the condition that $\phi$ is equivariant for the actions of $c$, where $z\in c$ acts on the source via $x\mapsto z\triangleright x$ and the action of $z\in c$ on the target is given by sending $\alpha(\bullet)\mapsto z\triangleright\alpha((z\triangleright)^{-1}\bullet)$, for $\alpha\in\mathrm{Hom}(c,c)$. Then, § 4.1(1) is the condition that these endomorphisms $\phi_{x}$ are automorphisms.
@@ 4.4. Decomposition of the comparison map
If $c^{\prime}$ is empty, then we claim that both the source and target of the map (4.1) are 00. We claim that some power of $U$ acts by 00 on the source and target, since a large enough power of $U$, $U^{j}$, is in the two sided ideal of $\pi_{0}A$ generated by $\alpha_{\beta}^{\operatorname{ord}(\beta)n_{\beta}}$. This is because by the pigeonhole principle, one can see that any element of $\pi_{0}\operatorname{Hur}^{G,c}_{n}$, written as a product of elements in $c$ for $n$ large enough contains $\max_{\beta\in c}(\operatorname{ord}(\beta)n_{\beta})$ many copies of a single $\beta\in c$. By pulling these to the left using the braid group action, we see the element is divisible by $\alpha_{\beta}^{\operatorname{ord}_{\beta}n_{\beta}}$. Since $U^{j}$ is in the ideal generated by the smallest powers of $\alpha_{\beta}^{\operatorname{ord}(\beta)}$ that act trivially, and hence it acts trivially. Since $U$ acts both as an isomorphism and nilpotently, the source and target of (4.1) with $c^{\prime}=\emptyset$ are both 00.
@@ 4.5. Showing the map is an equivalence
In this section we show that the maps in § 4.4 are an equivalence for each subrack $c^{\prime}\subset c$. In the case that $c^{\prime}=c$, the fact that the map in § 4.4 is an equivalence comes from the fact that the homology of the group completion of $\operatorname{Hur}^{c}$ can be identified with the rack homology of the rack $c$, which is computed in [EG03]. We verify this in § 4.5. In general, we show that completion at $\alpha_{c-c^{\prime}}$ (see § 4.4) corresponds to projecting in pointed spaces to the components of $\operatorname{Hur}^{c^{\prime}}$ at the level of homology, allowing us to again apply § 4.5 for the rack $c^{\prime}$ in this case. Identifying the completion in terms of $\operatorname{Hur}^{c^{\prime}}$ is at the core of our proof, and is proven using a homological epimorphism argument whose main input is a topological one identifying two bar constructions in pointed spaces as homotopy equivalent.
@@
§ 4.5 was essentially already known, and an alternate proof in the case that $c$ is a quandle with a single orbit under the action on itself proof can be found in [Bia24, §6.5].  [^5] We note that the argument presented in [Bia24, §6.5] describing the homology of the group completion isn’t quite correct, because it is claimed there that a certain algebra $\mathcal{A}(\mathcal{Q})$ is given by $\mathbb{Q}[x_{S}|S\in\operatorname{conj}(\mathcal{Q}_{+})]/(x^{2}_{S})$, but it should really be $\mathbb{Q}[x_{S}|S\in\operatorname{conj}(\mathcal{Q}_{+})]/(x_{S})^{2}$. In other words, the algebra written there is the homology of a product of $2$-spheres, while it should be the homology of a wedge of $2$-spheres. One can fairly easily correct this error and obtain a correct statement. We note that in the case there is only a single conjugacy class in $\mathcal{Q}_{+}$, which is the case relevant to the present article, the argument given in [Bia24, §6.5] is correct.
The proof we give here is modeled off of the argument in [RW20, Corollary 5.4]. We note that the argument given there has a small gap, where it is claimed on the first line of [RW20, p. 15] that $B\operatorname{Hur}^{c}_{G}\to B\operatorname{Conf}$ being a rational homology equivalence implies $\Omega B\operatorname{Hur}^{c}_{G}\to\Omega B\operatorname{Conf}$ is a rational homology equivalence on each path component. In general this does not hold, but it does hold when the spaces are weakly simple (see Section 4.5 below), as was very helpfully pointed out to us by Oscar Randal-Williams. This observation allows us to fix the small gap.
Definition 4.5.0.
A connected space $X$ is weakly simple if $\pi_{1}X$ acts trivially on $\pi_{n}X$ for $n>1$.
Lemma 4.5.0.
Suppose that $k$ is a localization of the integers and $X\to Y$ is a $k$-homology equivalence of weakly simple spaces that is also a $\pi_{1}$-equivalence. Then $\Omega X\to\Omega Y$ is also a $k$-homology equivalence.
@@
$\Omega X\to\Omega Y$ can be identified with the map $\pi_{1}X\times\Omega\tilde{X}\to\pi_{1}X\times\Omega\tilde{Y}$, where $\tilde{X},\tilde{Y}$ are the universal covers of $X,Y$ respectively. For simply connected spaces, $k$-homology equivalences and $k$-homotopy equivalences are the same, so it is enough to show that $\tilde{X}\to\tilde{Y}$ is a $k$-homotopy equivalence. We will show inductively on $i\geq 1$ that it is a $k$-homotopy equivalence in degrees $\leq i$. The base case $i=1$ holds by assumption. For the inductive step, if we know the result in degrees $\leq i$, then we aim to show it holds in degree $i+1$. Consider the fiber sequences $\tau_{\geq i+1}X\to X\to\tau_{\leq i}X$ and $\tau_{\geq i+1}Y\to Y\to\tau_{\leq i}Y$. The $k$-homology of $\tau_{\leq i}X$ and $\tau_{\leq i}Y$ agree by the inductive hypothesis. By the Hurewicz theorem and our assumption that $X,Y$ are weakly simple, the action of $\pi_{1}X$ on $H_{i+1}(\tau_{\geq i+1}X;k)=\pi_{i+1}X\otimes k$ and the action of $\pi_{1}Y$ on $H_{i+1}(\tau_{\geq i+1}Y;k)=\pi_{i+1}Y\otimes k$ are trivial. It follows by examination of the map of Serre spectral sequences of these fiber sequences and the fact that $X\to Y$ is a $k$-homology equivalence that the map $H_{i+1}(\tau_{\geq i+1}X;k)\to H_{i+1}(\tau_{\geq i+1}Y;k)$ is an equivalence. Using the Hurewicz theorem to identify $\pi_{i+1}(\tilde{X})\otimes k\simeq H_{i+1}(\tau_{\geq i+1}X;k)$, and similarly for $Y$ in place of $X$, we have $\pi_{i+1}(\tilde{X})\otimes k\simeq H_{i+1}(\tau_{\geq i+1}X;k)\simeq H_{i+1}(\tau_{\geq i+1}Y;k)\simeq\pi_{i+1}(\tilde{Y})\otimes k$, completing the induction. ∎
The following lemma identifies the target of the comparison map after group completion:
Lemma 4.5.0.
The map
@@
is an equivalence.
Proof.
We will first show that Note that $\pi_{1}(\psi):\pi_{1}(\Omega B\operatorname{Conf}^{c})\to\pi_{1}(\Omega B\operatorname{Conf}\times_{\Omega B\pi_{0}\operatorname{Conf}}\Omega B\pi_{0}\operatorname{Conf}^{c})$ is an equivalence. Therefore, in order to show $\psi$ is an equivalence, it suffices to show $\psi$ is an equivalence on homology. By the group completion theorem,
@@
Hence, $\psi$ is an equivalence. ∎
The lemma below is the key ingredient in the proof of the proposition. Recall that $c/c$ is the set of components of the rack $c$, i.e the orbits of the action of $c$ on itself. There is a natural surjective map of racks $c\to c/c$, where $c/c$ is given the trivial rack structure.
Lemma 4.5.0.
For any finite rack $c$, the map
$B\operatorname{Hur}^{c}\to B\operatorname{Hur}^{c/c}$
induced by $c\to c/c$ is a $\mathbb{Z}[\frac{1}{|G^{0}_{c}|}]$-homology equivalence.
Proof.
This lemma follows as in the proof of [RW20, Corollary 5.4]. Using Section A.4 for $X=c$, $M=N=*_{+}$, we obtain an explicit topological model for $B\operatorname{Hur}^{c}$. By filtering this space by the number of labels on the interior, we get a filtration such that the associated cohomology spectral sequence is the cochain complex computing rack cohomology of $c$, see [EG03]. It follows from [EG03, Theorem 4.2] that the map $c\to*$ induces an isomorphism on rack cohomology with $\mathbb{Z}[\frac{1}{G^{0}_{c}}]$-coefficients, which completes the proof. ∎
Proof of Section 4.5.
By the group completion theorem, $v[\alpha_{c}^{-1}]$ can be identified with the map $\Omega B\operatorname{Hur}^{c}\to\Omega B\operatorname{Conf}^{c}$ on homology. It is then enough to show that the map $\Omega B\operatorname{Hur}^{c}\to\Omega B\operatorname{Conf}^{c}$ is a rational homology equivalence.
We next wish to reduce to showing $B\operatorname{Hur}^{c}\to B\operatorname{Conf}^{c}$ is a rational homology equivalence. It is shown in [FRS07, Proposition 5.2] that for any rack $X$, $BX$ is weakly simple. By Section 4.5, $B\operatorname{Conf}^{c}$ is also weakly simple, since the action of $\pi_{1}$ on higher homotopy groups factors through $\pi_{1}B\operatorname{Conf}$ acting on the higher homotopy groups of $B\operatorname{Conf}$, and $B\operatorname{Conf}$ is the classifying space of a rack.
Applying Section 4.5 to the map $B\operatorname{Hur}^{c}\to B\operatorname{Conf}^{c}$, we learn that if it is a a rational homology equivalence, then $\Omega B\operatorname{Hur}^{c}\to\Omega B\operatorname{Conf}^{c}$ is also a rational homology equivalence. Here, we are also using that the map $\pi_{1}B\operatorname{Hur}^{c}\to\pi_{1}B\operatorname{Conf}^{c}$ is an equivalence by construction of $\operatorname{Conf}^{c}$.
@@
It remains then to show that the map $B\operatorname{Hur}^{c}\to B\operatorname{Conf}$ is an equivalence on rational homology, but this follows from applying Section 4.5, since $c/c=*$ since $c$ is connected, and $\operatorname{Hur}^{*}$ is $\operatorname{Conf}$. ∎
@@
For $c$ a rack and $c^{\prime}\subset c$ a subset, we define the normalizer of $c$ in $c^{\prime}$ to be $N_{c}(c^{\prime}):=\{x\in c|x\triangleright c^{\prime}=c^{\prime}\}$. We say that a subrack $c^{\prime}\subset c$ is self-normalizing if $c^{\prime}=N_{c}(c^{\prime})$.
@@
So, it is enough to verify the map (4.7) is a homology equivalence, which follows from § 4.5 below, using that $c^{\prime}$ is self-normalization so $N_{c}(c^{\prime})=c^{\prime}$. ∎
@@
Let $c$ be a rack, $c^{\prime}\subset c$ be a subrack with normalizer $N_{c}(c^{\prime})$, and $X_{+}$ be $\pi_{0}\operatorname{Hur}^{c^{\prime}}[\alpha_{c^{\prime}}^{-1}]_{+}$, viewed as a $\operatorname{Hur}^{c^{\prime}}_{+}$-bimodule. Then the map
(4.8) $X_{+}\otimes_{\operatorname{Hur}^{c}_{+}}X_{+}\to X_{+}\otimes_{\operatorname{Hur}^{N_{c}(c^{\prime})}_{+}}X_{+}$
@@
The map of interest $X_{+}\otimes_{\operatorname{Hur}^{c}_{+}}X_{+}\to X_{+}\otimes_{\operatorname{Hur}^{N_{c}(c^{\prime})}_{+}}X_{+}$ has a section induced from the inclusion of racks $N_{c}(c^{\prime})\to c$. It thus suffices to show that this section induces a homology equivalence.
We use the notation $\overline{Q}^{*}_{\epsilon}[M,\operatorname{Hur}_{+}^{N_{c}(c^{\prime})},N]$, as defined in § A.4 and Theorem A.4.1, which we describe below. By Theorem A.4.1, in order to prove this section is a homology equivalence it suffices to show the inclusion
$\iota_{\epsilon}:\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}_{+}^{N_{c}(c^{\prime})},X_{+}]\to\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}_{+}^{c},X_{+}]$
@@
The space $\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{N_{c}(c^{\prime})}_{+},X_{+}]$ is defined similarly, except the the labelling function $l$ must land in $N_{c}(c^{\prime})\subset c$.
At this point, it may be useful to refer to the notion of two spaces being ind-weakly equivalent, defined in § A.4. Let $S^{c,c^{\prime}}_{\epsilon}$ denote the quotient of the inclusion $\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{N_{c}(c^{\prime})}_{+},X_{+}]\to\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$.
@@
Note that any point in $S^{c,c^{\prime}}_{\epsilon}$ that is not the base point can be represented by a point in $\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$ where some $l(p_{i})$ is not in $N_{c}(c^{\prime})$.
It will be easier to prove the desired homology equivalence at the level of the associated graded of a certain filtration, which we define next. We define the filtration $F_{\bullet}S^{c,c^{\prime}}_{\epsilon}$ on $S^{c,c^{\prime}}_{\epsilon}$, where, for $j\geq 0$, $F_{j}S^{c,c^{\prime}}_{\epsilon}$ is the subset of $S^{c,c^{\prime}}_{\epsilon}$ consisting of those points where one of $l(p_{k-j}),\ldots,l(p_{k})$ lies in $N_{c}(c^{\prime})-c$, together with the base point. We set $F_{-1}S^{c,c^{\prime}}_{\epsilon}$ to be the basepoint, and use $G_{j}S^{c,c^{\prime}}_{\epsilon}$ to denote the associated graded of the filtration $F_{j}S^{c,c^{\prime}}_{\epsilon}/F_{j-1}S^{c,c^{\prime}}_{\epsilon}$, obtained by identifying $F_{j-1}S^{c,c^{\prime}}_{\epsilon}$ with the base point.
@@
We may choose a function $\xi:c-N_{c}(c^{\prime})\to c^{\prime}$ such that $g\triangleright\xi(g)\notin c^{\prime}$ for each $g\in c-N_{c}(c^{\prime})$, since $g\triangleright c^{\prime}\neq c^{\prime}$ from the definition of the normalizer of $c^{\prime}$.
@@
Suppose $y$ as above lies in $F_{j}Q_{\epsilon}^{c,c^{\prime}}$. Define $v$ to be the largest number so that $l(p_{v})\notin N_{c}(c^{\prime})$. By definition of the filtration $F_{j}Q_{\epsilon}^{c,c^{\prime}}$, we have $v\geq k-j$. Let $\epsilon^{\prime}:=\frac{\epsilon}{2}$. We now construct a continuous homotopy $H:F_{j}Q^{c,c^{\prime}}_{\epsilon}\times I\to G_{j}S^{c,c^{\prime}}_{\epsilon^{\prime}}$ as follows: If either $m,n$ is the base point, the point $y$ is equivalent to the base point under relation $(iii)$, and we choose the constant homotopy at the base point. So now assume this is not the case. If $v>k-j$, then the point is in $F_{j-1}Q_{\epsilon}^{c,c^{\prime}}$, so goes to the base point in the associated graded $G_{j}Q_{\epsilon}^{c,c^{\prime}}$. Thus, we may choose the constant homotopy on such points as well. So now we may assume as well that $v=k-j$.